Patient # Test A Test B Test C Disease Status

1

D D D D
2 D D D D
3 D D D D
4 D D D D
5 D D D D
6 D D D D
7 D D D D
8 D D D D
9 D D D D
10 D D D D
11 D D D D
12 D D D D
13 D D D D
14 D D D D
15 D D D D
16 D D D D
17 D D D D
18 D D D D
19 D D D D
20 D D D D
21 D D D D
22 D D D D
23 D D D D
24 ND D D D
25 ND D D D
26 ND D D D
27 ND D D D
28 D ND D D
29 D ND D D
30 D ND D D
31 D D ND D
32 ND ND D D
33 ND ND D D
34 ND ND D D
35 ND ND D D
36 ND ND D D
37 ND ND ND ND

Task is to calculate sensitivity and specificity for each of the tests.

D = Disease ND = No Disease

Question 1

If you were most concerned about Sensitivity, which is the best screening test?When you might be most concerned about Sensitivity? If you were most concerned about Specificity, which is the best screening test?

Test A ND Test+ Specificity Test B Specificity Test C Specificity
Answer #1

ANSWER

Test A

          D = 27, ND = 10

          Sensitivity.

                Formula = (True positives (TP) / (True positive(TP) + False negative(FN) ) * 100.

                True positives = 27

                False negative = 9

                         Sensitivity = (27 / (27 + 9)) 100

                                           = 0.75 * 100

                                           = 75%

          Specificity

             Formula =(True negatives(TN) / False positives(FP) + True negatives(TN)) * 100

                 True negative = 1

                 False positive = 0

                           Specificity = (1 / 0+1) 100

                                             = 1 * 100

                                             = 100%

Test B

          D = 28, ND = 9

         Sensitivity = (TP /TP+FN) * 100

                    TP = 28

                    FN = 8

                           = (28 / 28 + 8 ) 100

                           = 0.78* 100

                           = 78%.

         Specificity = (TN / FP + TN) 100

                       TN = 1

                        FP = 0

                          = (1 / 0+1) 100

                          = 1* 100

                          = 100%

Test C

        D = 35, ND = 2

       Sensitivity =( TP /TP + FN ) 100

                      TP = 35

                      FN = 1

                       = (35 /35 + 1) 100

                       = 0.97.2 * 100

                       = 97.2%

    Specificity = ( TN / FP + TN) 100

               TN = 1

               FP = 0

                    = (1 / 0+1) 100

                    = 100%

* Best screening test for sensitivity is Test C which giv4e more accurate information about

   positive rate.

* More concern about sensitivity is because it help in probability of detection or measures

percentage of people who identified with disease.

* All three test are best for measure specificity or true negative rate.